OpenStudy (anonymous):

For these functions, determine which ones are neither even nor odd a. f(x) = 3x^2 + |x| b. f(x) = 3x^3 + x^2 +x - 4 c. f(x) = 1/2x d. f(x) = 2x^5 - x^3 + 2x e. f(x) = 4x + 1/x f. f(x) = 2x^4 + x^2 - x + 2 I have to enter each letter but I am soo lost :(

OpenStudy (campbell_st):

the tests you need to replace x with -x if f(-x) = f(x) then the function is even e.g. f(x) = x^2 then f(-x) = (-x)^2 so f(-x) = x^2 hence f(-x) = f(x) if f(-x) = -f(x) then the function is odd e.g. f(x) = x^3 so f(-x) = (-x)^3 = -x^3 so f(-x) = - f(x) is its not odd nor even then its neither eg f(x) = 3x + 1 f(-x) = 3(-x) + 1 = -3x + 1 which is neither f(x) or - f(x) so its neither. so substitute -x into each and then check

OpenStudy (anonymous):

Nicely explained....

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Also,the graph of even function is symmetric about y-axis....

OpenStudy (campbell_st):

and as a hint, (a) is even